OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned?

OpenStudy (anonymous):

74

OpenStudy (anonymous):

74 men auditioned: 2x-3(120-x)=20 x=74 X= number of men (120-x)= number of women

sam (.sam.):

x represents the number of men. y represents the number of women, x + y = 120 2x - 3y = 20 Multiply the first equation by 3, so 3x + 3y = 360 2x - 3y = 20 Therefore 5x = 380 x = 380/5 = 76 y = 120 - 76 = 44

OpenStudy (anonymous):

which one 76 or 44 ?

sam (.sam.):

x represents the number of men. 76

OpenStudy (anonymous):

If twice the number of men less three times the number of women is equal to 20 m + w = 120 2m - 3w = 20 -> m = 76, w = 44